OpenStudy (anonymous):

He uses 0.24kg of wheat flour to make each loaf of bread. If he uses a combined total of 3.06kg of flour to make 6 loaves, how much corn flour is used in each loaf

OpenStudy (anonymous):

\[\Huge \color{Lime}{~\:\:\:\:\:\:\:\:\:\:\:\:\frak Welcome\:\:\frak To\:\:\frak Open\frak Study}\]

OpenStudy (anonymous):

let w = amount of wheat flour per loaf = 0.24 c = amount of corn flour per loaf then \[6(w)+6(c)=3.06\\ 6(0.24)+6c=3.06\\ 1.44+6c=3.06 \] continue and solve for "c"

OpenStudy (anonymous):

do i add 3.06 and 0.24?

OpenStudy (anonymous):

or 1.44 + 0.24

OpenStudy (anonymous):

6.24 *

OpenStudy (anonymous):

\[ 1.44+6c=3.06\\ -1.44\qquad-1.44\\ \;\\ \qquad6c=1.62\\ \text{now, divide both sides by 6} \]